# UniTS: A Unified Multi-Task Time Series Model

### v2

@@ Abstract
Advances in time series models are driving a shift from conventional deep learning methods to pre-trained foundational models. While pre-trained transformers and reprogrammed text-based LLMs report state-of-the-art results, the best-performing architectures vary significantly across tasks, and models often have limited scope, such as focusing only on time series forecasting. Models that unify predictive and generative time series tasks under a single framework remain challenging to achieve. We introduce UniTS, a multi-task time series model that uses task tokenization to express predictive and generative tasks within a single model. UniTS leverages a modified transformer block designed to obtain universal time series representations. This design induces transferability from a heterogeneous, multi-domain pre-training dataset—often with diverse dynamic patterns, sampling rates, and temporal scales—to many downstream datasets, which can also be diverse in task specifications and data domains. Across 38 datasets spanning human activity sensors, healthcare, engineering, and finance domains, UniTS model performs favorably against 12 forecasting models, 20 classification models, 18 anomaly detection models, and 16 imputation models, including repurposed text-based LLMs. UniTS demonstrates effective few-shot and prompt learning capabilities when evaluated on new data domains and tasks. In the conventional single-task setting, UniTS outperforms strong task-specialized time series models. The source code and datasets are available at https://github.com/mims-harvard/UniTS.
@@ 1 Introduction
Foundation models, especially large language models (LLMs), have transformed deep learning. Instead of training many task-specific models, a single pretrained model can support multiple tasks. Such models in language [9, 98] and vision [90, 49, 32] can be efficiently adapted to new tasks with little to no additional training via multi-task learning [122], few-shot learning [105, 84], and prompting [65]. Beyond vision and language, there is a need for similarly versatile time series models that can support time series from many domains—including medicine [34], engineering [99], and science [47]—and a range of tasks—forecasting, classification, imputation, and anomaly detection.
@@
Achieving multi-task time series models that unify predictive and generative tasks under a single model remains challenging. Inherently multi-domain time series datasets have varied temporal scales, sampling rates, and dynamic patterns, making them difficult to handle [121]. Within one model, existing time series models either cannot handle samples with different number of variables [109, 66, 14] or treat variables as independent ones [80]. Additionally, common time series tasks are diverse, with fundamentally different objectives and diverging task specifications across generative and predictive tasks. For example, generative task forecasting entails generating future values in a time series, while predictive task classification involves making discrete predictions, often made for an entire sample. Further, the same task across different datasets may require different specifications, such as generative tasks that vary in length and predictive tasks featuring multiple classes. Consequently, time series models remain largely task-specific, with unique architectures designed and trained from scratch for either forecasting [66, 80, 116], classification [30, 110], or other tasks [113, 109]. Recent attempts to pretrain unified models [36, 22] or reprogram LLMs to time series [115, 12, 126, 46, 95] still heavily rely on either extensive finetuning or adding and training task- and dataset-specific modules. A recent line of research explores the use of generative pre-training transformers for time series forecasting tasks [10, 115, 46, 28]; while these models report remarkable performance, they focus solely on forecasting tasks without considering other time series tasks. Each approach either burdens users with the need to design new modules and retrain models for each task or limits them to one type of time series task. Similar to how foundation models for language and vision define a wide class of tasks through unified space, a unified time series model must also unify task specifications that are either generative or predictive in nature. A unified model seeks to use one set of weights for multiple tasks, eliminating the need to train task-specific models from scratch. This approach enhances resource efficiency by supporting both generative and predictive tasks and rapid adaptation to new tasks and datasets.
@@
Figure 1: UniTS is a unified multi-task time series model that can process predictive and generative tasks across time series domains.
@@
Present work. To address these challenges, we introduce UniTS, a unified multi-task time series model that handles a broad range of time series tasks. We conduct an extensive comparison of UniTS to 12 forecasting methods, 20 classification methods, 18 anomaly detection methods, and 16 imputation methods, including transformer-based, LLM-based, RNN-based, and classic methods to demonstrate strong performance of UniTS and its generalizability to new tasks. This is accomplished through the following model design: 1) Task tokenization: UniTS converts task specifications into a unified token representation, enabling universal task specification without post hoc modifications to the network architecture. 2) Unified time series architecture: UniTS processes heterogeneous time series with diverse variables and lengths without modifying the network architecture. To achieve this, UniTS employs self-attention across both time and variable dimensions to adapt to varying temporal data dynamics. We introduce a dynamic linear operator to model dense relations between data points in the time dimension and a module to mitigate interference in the feature space of heterogeneous data. 3) Both generative and predictive tasks: The universal task specification and unified time series architecture enable UniTS to share weights across tasks through co-training on a wide variety of datasets. We adopt a masked reconstruction pretraining to jointly co-train the model for generative and predictive tasks.
@@
In the single-task setting, where models are trained separately for each dataset, UniTS outperforms strong task-specialized time series models and repurposed LLMs in forecasting, classification, anomaly detection, and imputation tasks. In a challenging multi-domain/task setting, we find that a single shared-weight UniTS model can successfully handle 38 tasks, demonstrating its potential as a multi-task time series model. UniTS outperforms leading baselines, which require data- and task-specific modules, achieving the highest average performance among all tasks and excelling in 27 out of 38 tasks. Additionally, UniTS can perform prompt-based learning and zero-shot direct multi-step forecasting with new lengths, which existing models that use task- and data-specific heads do not support. In direct multi-step forecasting with new lengths, UniTS outperforms the strongest baseline (which relies on sliding windows) by 10.5%. UniTS can be adapted to new tasks via parameter-efficient prompt learning, achieving performance comparable to UniTS’s fully fine-tuned counterpart. Across 20 forecasting datasets, prompted UniTS outperforms the fully fine-tuned version, reducing MAE from 0.381 to 0.376. UniTS also exhibits few-shot transfer learning, effectively handling tasks such as imputation, anomaly detection, and out-of-domain forecasting and classification without requiring specialized data- or task-specific modules. For instance, UniTS outperforms the strongest baseline by 12.4% (MSE) on imputation tasks and 2.3% (F1-score) on anomaly detection tasks. UniTS paves the way towards more comprehensive and unified time series models. The source code and datasets are available at https://github.com/mims-harvard/UniTS.
@@ 2 Related Work
Traditional time series modeling. Time series analysis has been extensively explored in both the statistics and machine learning communities for many years [44, 100, 120, 18, 78]. Numerous neural architectures have been developed for specific time series tasks such as forecasting [111, 64, 67, 66, 104], classification [112, 70, 69], anomaly detection [25, 55, 16], and imputation [17, 48, 3]. Task-specific models are typically trained via supervised learning on individual datasets, necessitating specialized modules. For example, a classification model requires a classification head with a specific number of classes, while data processing modules must handle a predetermined number of variables. In contrast, UniTS aims to unify various tasks into a universal task specification, enabling the handling of diverse data with a single, unified network architecture. This approach facilitates training a multi-task model capable of addressing multiple time series tasks simultaneously.
@@
General time series modeling. Foundation models are trained on broad data at scale to support diverse tasks with no or minimal additional training [8]. They have reshaped research areas, including language modeling [9, 98] and computer vision [61, 49]. Recent studies in time series analysis have sought to develop models with similar capabilities. This includes developing novel architectures to capture diverse time series signals. For instance, TimesNet [109] uses multiple levels of frequency-based features obtained through Fourier transform to capture complex time series signals. Additionally, there have been efforts to reprogram LLMs for time series tasks [79, 12, 126, 46, 10]. GPT4TS [126] and Time-LLM [46], fine-tune embedding layers or align time series sample with the LLM-based text prototypes (GPT-2 [87]) for time series tasks. Lag-Llama [88] pretrains a model on time series data from several domains for forecasting tasks. Concurrently, the Moment model [36] is pretrained on a diverse set of time series data. However, these approaches still require task-specific modules and tuning for each task. In contrast, our UniTS model supports both generative and predictive tasks without the need for extensive task-specific model adjustments.
@@
Prompt learning. Prompt learning has emerged as an efficient method for task adaptation in large models [54, 86, 118, 13, 41]. Some approaches construct prompts directly in the model’s input domain, such as text prompts for LLMs [2]. Other methods involve tuning soft token inputs to frozen language models [57]. In time series, PromptCast [115] and LLMTime [79] converts time series data into prompts for LLMs to facilitate forecasting. TEMPO [10] is another prompt-based approach that uses a learned set of prompts for LLM-based forecasting applications, while GPT4MTS [45] integrates both textual and numerical data to fine-tune LLMs for forecasting. In contrast, UniTS is trained exclusively on time series data, eliminating the need for computationally expensive pre-trained LLMs. Moreover, the universal task tokenization enables a frozen UniTS to adapt to new tasks beyond forecasting, such as classification and imputation. Further discussion of related work can be found in Appendix A.
@@ 3 Problem Formulation
Notation. We are given a set of multi-domain datasets $\mathcal{D}=\{\mathcal{D}_{i}|i=1,\ldots,n\}$, where each dataset $\mathcal{D}_{i}$ can have a varying number of time series samples; samples can be of varying time lengths and have varying numbers of sensors/variables. Each dataset is described as $\mathcal{D}_{i}=(\mathcal{X}_{i},\mathcal{Y}_{i})$, where $\mathcal{X}_{i}$ denotes time series samples and $\mathcal{Y}_{i}$ specifies a task defined on $\mathcal{X}_{i}$. Let $\mathcal{X}$ and $\mathcal{Y}$ be collections, defined as $\mathcal{X}=\{\mathcal{X}_{i}|i=1,\ldots,n\}$ and $\mathcal{Y}=\{\mathcal{Y}_{i}|i=1,\ldots,n\}$, respectively. A time series sample in datasets is denoted as $\mathbf{x}\in\mathbb{R}^{t\times v}$, where $t$ and $v$ are the length of time series sample and the number of variables, respectively. We use time dimension and variable dimension to indicate the row and column dimensions in $\mathbf{x}$. $\mathcal{Y}_{i}$ contains four common time series tasks: forecasting, classification, anomaly detection, and imputation. Further, each task type can be instantiated in numerous ways, e.g., forecasting over different time lengths and classification with varying numbers of classes. We use $F(\mathcal{X},\theta)$ to denote a multi-task model trained on $\mathcal{X}$. See Table 12 for notation details.
@@
To realize the above desiderata, UniTS supports multi-task, prompt-based, zero-shot, and few-shot learning. Multi-task learning: UniTS specifies a single model $F(\mathcal{X},\theta)\rightarrow\mathcal{Y}$ for tasks $\mathcal{Y}$ defined on datasets $\mathcal{X}$. Multi-task learning showcases the flexibility of the model to learn across time series domains and tasks. Prompt learning: By leveraging prompt tokens, UniTS supports prompt learning, $\textit{Prompting}\{F(\mathcal{X},\theta),\text{token}\}\rightarrow\mathcal{Y}$, across tasks while keeping the model frozen. Additionally, UniTS can be trained in a single-task manner, following the same setup as used by many existing models. Other settings are described in Appendix C.1.
@@ 4 UniTS Model
UniTS is a multi-task model with a unified network architecture. It uses a token-based format to describe tasks and time series from different domains. We introduce a novel approach with three distinct token types: sample, prompt, and task tokens, each serving a unique purpose in time series analysis. The input time series sample is tokenized into sample tokens. Prompt tokens provide essential context for the task, guiding the model to accomplish the user-specified task. Task tokens (GEN and CLS) are combined with other tokens and used for generative and predictive tasks. UniTS then processes task tokens by converting them into task predictions to produce the final model output. Unlike tokens in transformers such as PatchTST [80], UniTS introduces new features for time series analysis: sample tokens support multivariate time series instead of univariate data, prompt tokens enable efficient multi-task and prompt learning for time series analysis, similar to LLMs [98], and task tokens unify predictive and generative tasks into one format.
@@ 4.1 Prompting UniTS with Unified Time Series Data Tokens
Sample tokens. We divide time series input sample $\mathbf{x}\in\mathbb{R}^{t\times v}$ into patches along the time dimension using a non-overlapping patch size of $k$. A linear layer projects each patch into an embedding vector of length $d$, obtaining sample tokens $\mathbf{z}_{\mathbf{x}}\in\mathbb{R}^{s\times v\times d}$, where $s=t/k$. Since $v$ and $s$ varies across time series data domains, we keep the variable and time dimension in tokens. $\mathbf{z}_{\mathbf{x}}$ are then added with learnable positional embeddings. Prompt tokens. Prompt tokens $\mathbf{z}_{p}\in\mathbb{R}^{p\times v\times d}$ are defined as learnable embeddings, where $p$ is the number of tokens. In a multi-task setting, each dataset has its own set of prompt tokens. These tokens incorporate the specific context related to the data and the task that the model needs to complete. For prompt learning, with the pre-trained model weights being frozen, UniTS adapts to new tasks by utilizing prompt tokens learned with the prompt tuning. Prompt learning is more efficient than tuning new data/task-specific heads and achieves comparable performance to full model finetuning, as shown by few-shot learning experiments on new tasks (Tables 4 and 5) and new datasets (Table 3).
@@
which is then fed into the model. We define class embeddings $\mathbf{z}_{e}\in\mathbb{R}^{e\times v\times d}$ for each of $e$ classes in the task. These class embeddings are either trained or generated by averaging CLS tokens of training samples in each class. Finally, the class for sample $\mathbf{x}$ is predicted by finding the class embedding vector in $\mathbf{z}_{e}$ that is the closest to the CLS token $\mathbf{z}_{c}$:
@@
For imputation, missing values are imputed using the GEN tokens. For anomaly detection, the model takes a time series sample that can contain any number of potentially anomalous values, generates the output sample by reading-out the sample tokens, and then determines anomalous values based on the reconstruction error between the input sample and the generated sample. Details on using tokens for imputation and anomaly detection tasks are in Appendix C.2. All tokens and embeddings are trained to achieve their functions.

### v3

@@ Abstract
Although pre-trained transformers and reprogrammed text-based LLMs have shown strong performance on time series tasks, the best-performing architectures vary widely across tasks, with most models narrowly focused on specific areas, such as time series forecasting. Unifying predictive and generative time series tasks within a single model remains challenging. We introduce UniTS, a unified multi-task time series model that utilizes task tokenization to integrate predictive and generative tasks into a single framework. UniTS employs a modified transformer block to capture universal time series representations, enabling transferability from a heterogeneous, multi-domain pre-training dataset—characterized by diverse dynamic patterns, sampling rates, and temporal scales—to a wide range of downstream datasets with varied task specifications and data domains. Tested on 38 datasets across human activity sensors, healthcare, engineering, and finance, UniTS achieves superior performance compared to 12 forecasting models, 20 classification models, 18 anomaly detection models, and 16 imputation models, including adapted text-based LLMs. UniTS also demonstrates strong few-shot and prompt capabilities when applied to new domains and tasks. In single-task settings, UniTS outperforms competitive task-specialized time series models. Code and datasets are available at https://github.com/mims-harvard/UniTS.
@@ 1 Introduction
Foundation models, particularly large language models (LLMs), have transformed deep learning by enabling a single pre-trained model to support multiple tasks, eliminating the need for task-specific models. Language and vision models [9, 101, 92, 50, 32] can be adapted to new tasks with minimal additional training through approaches such as multi-task learning [125], few-shot learning [108, 86], and prompting [66]. Beyond language and vision, there is a growing need for similarly versatile models in time series that can accommodate data from diverse domains—including medicine [34], engineering [102], and science [48]—and support a wide range of tasks, such as forecasting, classification, imputation, and anomaly detection.
@@
Developing multi-task time series models that unify predictive and generative tasks under a single framework remains an open challenge. Time series datasets span multiple domains and exhibit varied temporal scales, sampling rates, and dynamic patterns, making them complex to manage [124, 78]. Existing models often fall short in adaptability, as they either struggle to handle samples with varying numbers of variables [112, 67, 14] or treat each variable as independent, overlooking important interdependencies [82]. Time series tasks are also highly diverse, encompassing distinct objectives and specifications across generative and predictive tasks. For example, generative forecasting tasks aim to produce future values within a time series, while predictive tasks may involve making discrete predictions for entire samples. Additionally, task requirements can vary significantly even within the same task type; for instance, generative tasks may involve different forecast lengths, and predictive tasks may feature multiple classification categories. As a result, time series models have mainly remained task-specific, with unique architectures typically designed and trained from scratch for forecasting [67, 82, 119], classification [30, 113], or other specialized tasks [116, 112]. Recent efforts to pre-train unified models [36, 22] or adapt LLMs for time series [118, 12, 129, 47, 97, 100] still heavily depend on extensive fine-tuning or the addition of task- and dataset-specific modules. Some models have explored generative pre-training transformers specifically for time series forecasting [10, 118, 47, 28], reporting strong results but focusing exclusively on forecasting without addressing other types of time series tasks. Consequently, these approaches require users to design and train new modules for each task or limit their application to a single type of tasks. To achieve a versatile, unified time series model—akin to foundational models in vision and language that operate across unified task spaces—a model must accommodate both generative and predictive tasks. Such a unified model would leverage a single set of weights for multiple tasks, removing the need to develop task-specific models from scratch. This approach would support a broad range of tasks and facilitate rapid adaptation to new datasets.
@@
Figure 1: UniTS is a unified multi-task time series model for predictive and generative tasks.
@@
Present work. To address these challenges, we introduce UniTS, a unified multi-task time series model capable of handling a broad spectrum of time series tasks. We rigorously compare UniTS against 12 forecasting methods, 20 classification methods, 18 anomaly detection methods, and 16 imputation methods, including transformer-based, LLM-based, RNN-based, and traditional approaches, to highlight UniTS’s generalizability to new tasks. This capability is achieved through the following model design: 1) Task tokenization: UniTS encodes task specifications into a unified token representation, enabling universal task specification without post-hoc architectural modifications. 2) Unified time series architecture: UniTS processes heterogeneous time series data with varying numbers of variables and sequence lengths without altering its network structure. To accomplish this, UniTS employs self-attention across time and variable dimensions to adapt to diverse temporal dynamics. We introduce a dynamic linear operator to model complex relationships between data points along the time dimension and a module to reduce interference in the feature space of heterogeneous data. 3) Support for generative and predictive tasks: The combination of universal task specification and a unified time series architecture allows UniTS to share weights across tasks by co-training on multiple datasets. We use a masked reconstruction pre-training approach, enabling UniTS to be jointly optimized for generative and predictive tasks.
@@
In the single-task setting, where models are trained individually for each dataset, UniTS outperforms task-specialized time series models and repurposed LLMs across forecasting, classification, anomaly detection, and imputation. In a challenging multi-domain, multi-task setting, we find that a single shared-weight UniTS model successfully handles 38 tasks, demonstrating its versatility as a multi-task time series model. UniTS surpasses top baselines that rely on data- and task-specific modules, achieving the highest average performance across tasks and excelling in 27 out of 38 tasks. Additionally, UniTS supports prompt-based learning and direct multi-step forecasting with flexible sequence lengths, capabilities not offered by models using task- and data-specific heads. In direct multi-step forecasting, UniTS outperforms the strongest baseline (which uses a sliding-window approach) by 10.5%. UniTS can also adapt to new tasks through parameter-efficient prompting, achieving results comparable to its fully fine-tuned counterpart. For example, across 20 forecasting datasets, prompted UniTS slightly outperforms the fully fine-tuned model, reducing MAE from 0.381 to 0.376. Furthermore, UniTS demonstrates effective few-shot transfer, successfully addressing tasks like imputation, anomaly detection, and out-of-domain forecasting and classification without requiring specialized modules. For instance, UniTS improves on the strongest baseline by 12.4% in MSE on imputation and 2.3% in F1-score on anomaly detection. UniTS paves the way toward unified time series models, offering strong performance and adaptability across tasks and domains.
@@ 2 Related Work
Traditional time series modeling. Time series analysis has been extensively explored in both the statistics and machine learning communities for many years [45, 103, 123, 18, 80]. Numerous neural architectures have been developed for specific time series tasks such as forecasting [114, 65, 68, 67, 107], classification [115, 71, 70], anomaly detection [25, 56, 16], and imputation [17, 49, 3]. Task-specific models are typically trained via supervised learning on individual datasets, necessitating specialized modules. For example, a classification model requires a classification head with a specific number of classes, while data processing modules must handle a predetermined number of variables. In contrast, UniTS aims to unify various tasks into a universal task specification, enabling the handling of diverse data with a single, unified network architecture. This approach facilitates training a multi-task model capable of addressing multiple time series tasks.
@@
General time series modeling. Foundation models, including language models [9, 101] and vision models [62, 50], are trained on broad data at scale to address diverse tasks with no or minimal additional training [8]. Recent studies in time series analysis have sought to develop models with similar capabilities. This includes developing novel architectures to capture diverse time series signals. For instance, TimesNet [112] uses multiple frequency-based features obtained through Fourier transform to capture complex time series signals. There have been several efforts to reprogram LLMs for time series tasks [81, 12, 129, 47, 10]. Models such as GPT4TS [129] and Time-LLM [47] adapt LLMs by fine-tuning their embedding layers or aligning time series samples with LLM-based text prototypes (e.g., GPT-2 [89]). Unlike these models, UniTS is trained exclusively on time series data rather than relying on LLM architectures. Another approach, Lag-Llama [90], pre-trains a model on time series data from multiple domains specifically for forecasting tasks. Similarly, the Moment model [36] is pre-trained on a diverse range of time series data. However, these approaches still require task-specific modules and tuning for each task. In contrast, our UniTS model supports generative and predictive tasks without requiring extensive task-specific model adjustments.
@@
Prompt learning. Prompt learning has emerged as an efficient method for task adaptation in large models [55, 88, 121, 13, 42]. Some approaches construct prompts directly in the model’s input domain, such as text prompts for LLMs [2]. Other methods involve tuning soft token inputs to frozen language models [58]. In time series, PromptCast [118] and LLMTime [81] convert time series data into prompts for LLMs to facilitate forecasting. TEMPO [10] is another prompt-based approach that uses a learned set of prompts for LLM-based forecasting applications, while GPT4MTS [46] integrates both textual and numerical data to fine-tune LLMs for forecasting. In contrast, UniTS is trained exclusively on time series data, eliminating the need for computationally expensive pre-trained LLMs. Moreover, the universal task tokenization enables a frozen UniTS to adapt to new tasks beyond forecasting, such as classification and imputation. Further discussion of related work can be found in Appendix A.
@@ 3 Problem Formulation
Notation. We are given a set of multi-domain datasets $\mathcal{D}=\{\mathcal{D}_{i}|i=1,\ldots,n\}$, where each dataset $\mathcal{D}_{i}$ can have a varying number of time series samples; samples can be of varying time lengths and have varying numbers of sensors/variables. Each dataset is described as $\mathcal{D}_{i}=(\mathcal{X}_{i},\mathcal{Y}_{i})$, where $\mathcal{X}_{i}$ denotes time series samples and $\mathcal{Y}_{i}$ specifies a task defined on $\mathcal{X}_{i}$. Let $\mathcal{X}$ and $\mathcal{Y}$ be collections, defined as $\mathcal{X}=\{\mathcal{X}_{i}|i=1,\ldots,n\}$ and $\mathcal{Y}=\{\mathcal{Y}_{i}|i=1,\ldots,n\}$, respectively. A time series sample in datasets is denoted as $\mathbf{x}\in\mathbb{R}^{t\times v}$, where $t$ and $v$ are the length of the time series sample and the number of variables, respectively. We use time dimension and variable dimension to indicate the row and column dimensions in $\mathbf{x}$. $\mathcal{Y}_{i}$ contains four common time series tasks: forecasting, classification, anomaly detection, and imputation. Further, each task type can be instantiated in numerous ways, e.g., forecasting over different time lengths and classification with varying numbers of classes. We use $F(\mathcal{X},\theta)$ to denote a multi-task model trained on $\mathcal{X}$. See Table 12 for notation details.
@@
To realize the above desiderata, UniTS supports multi-task, prompt-based, and few-shot learning. Multi-task learning: UniTS specifies a single model $F(\mathcal{X},\theta)\rightarrow\mathcal{Y}$ for tasks $\mathcal{Y}$ defined on datasets $\mathcal{X}$. Multi-task learning showcases the flexibility of the model to learn across time series domains and tasks. Prompt learning: By leveraging prompt tokens, UniTS supports prompt learning, $\textit{Prompting}\{F(\mathcal{X},\theta),\text{token}\}\rightarrow\mathcal{Y}$, across tasks while keeping the model frozen. Additionally, UniTS can be trained in a single-task manner, following the same setup as used by many existing models. Other settings are described in Appendix C.1.
@@ 4 UniTS Model
UniTS is a multi-task model with a unified network architecture. It uses a token-based format to describe tasks and time series from different domains. We introduce a novel approach with three distinct token types: sample, prompt, and task tokens, each serving a unique purpose in time series analysis. The input time series sample is tokenized into sample tokens. Prompt tokens provide essential context for the task, guiding the model to accomplish the user-specified task. Task tokens (GEN and CLS) are combined with other tokens and used for generative and predictive tasks. UniTS then converts task tokens into task predictions to produce the final model output. Unlike transformers such as PatchTST [82], UniTS introduces new token types: sample tokens allow for modeling of multivariate time series, prompt tokens enable efficient multi-task and prompt learning [101], and task tokens unify predictive and generative tasks into one format.
@@ 4.1 Prompting UniTS with Unified Time Series Data Tokens
Sample tokens. We divide time series input sample $\mathbf{x}\in\mathbb{R}^{t\times v}$ into patches along the time dimension using a non-overlapping patch size of $k$. A linear layer projects each patch into an embedding vector of length $d$, obtaining sample tokens $\mathbf{z}_{\mathbf{x}}\in\mathbb{R}^{s\times v\times d}$, where $s=t/k$. Since $v$ and $s$ vary across time series data domains, we keep the variable and time dimension in tokens. $\mathbf{z}_{\mathbf{x}}$ are then added with learnable positional embeddings.
Prompt tokens. Prompt tokens $\mathbf{z}_{p}\in\mathbb{R}^{p\times v\times d}$ are defined as learnable embeddings, where $p$ is the number of tokens. In a multi-task setting, each dataset has its own set of prompt tokens. These tokens incorporate the specific context related to the data and the task the model needs to complete. For each sample in the dataset, these prompt tokens are appended to the sample tokens and sent to the network to provide context information about the current sample. For prompt learning, with the pre-trained model weights being frozen, UniTS adapts to new tasks by utilizing prompt tokens learned with the prompt tuning. Prompt learning is more efficient than tuning new data/task-specific heads and achieves comparable performance to full model fine-tuning, as shown by few-shot learning experiments on new tasks (Tables 4 and 5) and new datasets (Table 3).
@@
which is then fed into the model. We define class embeddings $\mathbf{z}_{e}\in\mathbb{R}^{e\times v\times d}$ for each of $e$ classes in the task. These class embeddings are either trained or generated by averaging CLS tokens of training samples in each class. Finally, the class for sample $\mathbf{x}$ is predicted by finding the class embedding vector in $\mathbf{z}_{e}$ that is the closest to the CLS token $\mathbf{z}_{c}$ from the model output:
@@
For imputation, missing values are imputed using the GEN tokens. For anomaly detection, the model takes a time series sample containing any number of potentially anomalous values, generates the output sample by reading out the sample tokens, and then determines anomalous values based on the reconstruction error between the input sample and the generated sample. Details on using tokens for imputation and anomaly detection are in Appendix C.2. All tokens and embeddings are trained to achieve their functions.